# Membership Inference Attack with Partial Features

### v1

@@ Abstract
Machine learning models have been shown to be susceptible to membership inference attack, which can be used to determine whether a given sample appears in the training data. Existing membership inference methods commonly assume that the adversary has full access to the features of the target sample. This assumption, however, does not hold in many real-world scenarios where only partial features information is available, thereby limiting the applicability of these methods. In this work, we study an inference scenario where the adversary observes only partial features of each sample and aims to infer whether this observed subset was present in the training set of the target model. We define this problem as Partial Feature Membership Inference (PFMI). To address this problem, we propose MRAD (Memory-guided Reconstruction and Anomaly Detection), a two-stage attack framework. In the first stage, MRAD optimizes the unknown feature values to minimize the loss of the sample. In the second stage, it measures the deviation between the reconstructed sample and the training distribution using anomaly detection. Empirical results demonstrate that MRAD is effective across a range of datasets, and maintains compatibility with various off-the-shelf anomaly detection techniques. For example, on STL-10, our attack achieves an AUC of around 0.6 even with 40% of the missing features.
@@ I Introduction
Machine learning has been widely applied across various domains, such as medical diagnosis[1], traffic analysis[2], and autonomous driving[3]. However, current machine learning models remain vulnerable to privacy inference attacks. Adversaries can exploit interactions with the target model to infer privacy information from its training data, posing significant threats to user confidentiality and data privacy. A well-known example of such attacks is the Membership Inference Attack (MIA)[4, 5], which can be used to determine whether certain samples were included in the training set of a target model. This type of attack highlights the risk of information leakage during inference, where adversaries can extract sensitive training information at a low cost.
Membership inference attack has been shown to compromise the privacy of various machine learning models[6, 7, 8, 9]. In this paper, we focus on this attack against classification models. Existing membership inference attack can be roughly categorized into two types based on the attacker’s objective. The first type, exact‑match MIA, aims to determine whether a given sample exactly matches one of the samples in the training set of the target model, which is where most existing attacks fall[10, 11, 12]. These can be further divided into confidence‑based attacks[13, 14], label-only attacks[10, 15], and blind attacks[16]. The second type, range‑match MIA, does not seek an exact sample match. Instead, it aims to infer whether the model’s training set contains any sample with features in a specified range. In recent work, Tao et al.[17] were the first to formally define and investigate this setting, highlighting that membership inference can be extended beyond exact‑match assumptions.
@@
Inspired by Tao et al.[17], we pose a question: How much private information can membership inference attack extract when some features are missing? This setting aligns more closely with practical scenarios. For instance, consider an attacker attempting to determine whether an individual, Alice, has a particular disease by querying a diagnostic system deployed in a hospital she frequently visits. If Alice is indeed diagnosed with the disease, it is likely that her data is included in the training set. However, due to the difficulty of acquiring complete information, the attacker may only know a subset of her attributes, such as age, gender, and race. Applying a standard membership inference attack in this setting would require exhaustively enumerating all possible combinations of the unknown features, which is computationally infeasible and inefficient.
@@
As Figure 1 shows, in this paper, we consider a setting where the attacker has access to only a subset of a sample’s features and aims to infer whether this partial feature combination was present in the training set of a target model. We refer to this type of attack as Partial Feature Membership Inference (PFMI). In this setting, the observed partial features is termed known features, while the remaining unobserved ones are referred to as unknown features. The property existence attack, proposed as a generalization of membership inference [18], shares certain similarities with PFMI in that both seek to infer the presence of specific feature combinations. However, in property existence attack, the adversary can directly access complete samples containing the target feature combination [18, 19], whereas PFMI operates under stricter constraints by relying solely on partial observations. Although the attacker’s limited access to only partial feature information makes the inference task more challenging, this constraint also reflects a more realistic threat model. It shows that even with incomplete information, a resource-constrained adversary can still launch effective privacy inference attack. This amplifies the risks posed to current privacy protection mechanisms, which often assume attackers have access to complete samples.
@@
To tackle this problem, we propose MRAD—a two-stage attack framework that combines Memory-guided Reconstruction with Anomaly Detection. MRAD is specifically tailored to the partial feature setting and is capable of accurately inferring membership even when some input features are missing. Specifically, when we try to minimize the loss of the target sample by updating unknown features, reconstructed samples with non-member features often deviate significantly from the original distribution, resulting in anomalous data. We leverage anomaly detection techniques to identify such deviations, thereby determining whether the observed partial feature combination exists in the training set. Our main contributions can be summarized as follows.
@@
  • Attack Formalization under Partial Observability: We investigate a more realistic setting for membership inference attack, in which the adversary has access only to partial features of the target sample. We refer to this scenario as PFMI, and we provide a formal definition of the attack framework under this constraint.
@@
  • Two-Stage Attack Framework: We propose a two-stage attack framework designed for this new inference scenario. The framework is modular and can be integrated with a variety of off-the-shelf anomaly detection techniques.
@@
  • Extensive Evaluation: We extensively evaluate our attack framework on both image and tabular datasets, demonstrating its compatibility with various anomaly detection methods. In the STL-10 and Epsilon datasets, our method achieves an average AUC of 0.6 even with 40% missing features.
@@
  • In-depth Analysis & Case Study: We show how hyperparameter settings and the importance of observed features affect attack performance by experiments. We also present a case study to demonstrate how our attack can lead to privacy leakage in a real world scenario.
@@
The remainder of this paper is organized as follows. Section II reviews related work on membership inference and anomaly detection. Section III introduces the PFMI setting, including the attack objective, adversary capability and attack definition. Section IV presents the technical details of the proposed two-stage attack framework. Section V presents experimental results, including studies that validate design intuition, followed by a comprehensive evaluation of the attack performance in various settings. Section VI discusses potential defense mechanisms, analyzes the relationship between PFMI and other inference attacks, and outlines directions for future work. Section VII concludes the paper.
@@ II-A Membership Inference Attack
Membership Inference Attack Based on Exact Sample Matching: Membership inference attack against machine learning models was first introduced by Shokri et al.[4]. A binary classifier was trained as the attack model using prediction vectors produced by shadow models. To launch the attack, the target model’s prediction on a given input is provided to the attack model, which decides whether the input was included in the training data. Most follow-up studies adopt a similar setting, where the adversary has access to both the full prediction vector and the ground-truth label of the sample. These methods typically rely on the observation that member samples tend to yield lower loss values than non-members [13, 12, 5]. Song et al.[13] predict membership with a modified loss function. Ye et al.[14] proposed a membership inference attack that achieves an arbitrary false positive rate by using a non-member set to calibrate the decision threshold. Carlini et al. [5] trained multiple shadow models to examine the loss distributions of samples when they are in and out of the training set, further improving the attack performance under low false positive rates. Instead of using loss values for membership inference, Zarifzadeh et al.[11] computed a likelihood ratio based on the output confidences and used it as the inference metric.
@@
In the label-only scenario, the adversary can only observe the predicted label rather than the full confidence vector [15, 20]. Li et al. [15] leveraged the fact that member samples are often farther from the decision boundary, while Peng et al. [10] crafted minimally perturbed adversarial examples using a shadow model and inferred membership based on whether the perturbed samples were misclassified by the target model. Since their method relies solely on the final prediction of the model, it requires only a single query.
Another line of work, which is called blind membership inference, considers the setting where the adversary has access to the full prediction vector of the target sample but lacks knowledge of the true label. Hui et al. [16] proposed a blind black-box attack that performs differential comparisons between samples to infer membership without ground-truth guidance.
@@
Membership Inference Attack Based on Feature Range Matching: Tao et al.[17] proposed a novel form of membership inference attack that shifts the focus from individual samples to feature regions. Rather than determining whether a specific sample was included in the training set, their method aims to infer whether any data point within a given range appears in the training data. To perform the attack, they first sample instances from the specified range, apply standard membership inference techniques to assign scores to these samples, and then aggregate the scores to reach a final inference decision.
@@ II-B Anomaly Detection
Anomalous data refers to samples that deviate from the expected or learned data distribution. Two common categories of such anomalies are concept drift and out-of-distribution (OOD) samples. Concept drift occurs when the test-time data distribution shifts from the training distribution, leading to degraded model performance [21]. Yang et al. [22] detect such drift by modeling input data with autoencoders and computing Median Absolute Deviation(MAD) values between the current and original distributions. Wan et al. [23] proposed MCD-DD, which uses a dynamically updated encoder and Mahalanobis distance-based criteria to detect distributional changes in time series. OOD detection has also received significant attention, as the presence of OOD samples can bias model behavior. Park et al. [24] improved OOD detection by reducing overconfidence in OOD regions. Reiss et al. [25] designed a novel contrastive loss function that enhances contrastive learning for OOD detection. Liu et al. [26] exploited the clustering effect in the penultimate layer of neural networks to develop a post-hoc OOD detection method.
@@ III-A Attack Model
Attack Objective: In this work, the goal of the attacker is similar to that of traditional MIA, but with a key constraint: the attacker only has access to partial features of the target sample. The objective is to determine whether this particular feature combination exists in the training set of the target model. We refer to such feature combinations as member features if they appear in the training set, and non-member features otherwise. Apparently, any combination drawn from the training set is considered as member features.
Adversary Capability: We consider a white-box threat model in which the attacker has full access to the model, including its output, training algorithm, internal parameters, architecture, and gradients. Although the specific target sample may have missing features due to privacy concerns or incomplete collection, we assume that the attacker has knowledge of the training data distribution and can sample fully observed auxiliary data from it to support the attack, which is consistent with the prior inference attack settings[4, 13, 14, 11]. This is feasible in practice, as attackers may obtain partial data through social engineering or cyberattacks. However, the attacker is unable to obtain any complete sample that contains the exact feature combination to be inferred.
@@ III-B Attack Definition
Before introducing our proposed attack definition, we first revisit the standard definition of membership inference attack, as formalized in prior work[27, 11] using an indistinguishability game.
@@ 3. III
In this paper, for the sake of generality, we define partial features as a subset of the full feature vector, where the known features can appear at arbitrary positions. That is, the known and unkonwn features are not required to form a contiguous block or follow any fixed pattern. This flexible definition better captures real-world scenarios where feature availability is influenced by privacy constraints or unstable data sources.
@@ IV-A Design Intuition
In designing our attack, we build on a well‑established observation: overfitted models memorize their training data, causing training members to occupy regions of the input space associated with low loss values[4, 28, 29]. Specifically, for a supervised machine learning algorithm $\mathcal{T}$, the training process aims to learn a model that can effectively distinguish between samples from different classes. Since it is impractical to access the full data distribution, a finite subset of labeled examples $\mathcal{D}_{train}$, the training set, is typically sampled to represent it. A function $f_{\theta}$, parameterized by $\theta$, is then learned to map each input $x$ from $\mathcal{D}_{train}$ to its corresponding label $k$ by minimizing a loss function $\mathcal{L}_{\theta}(x,k)$. Because the optimization only considers the training set, the model tends to assign lower loss values to training samples compared to unseen ones. This results in loss valleys around training points in the input space, a phenomenon that forms the basis of many membership inference attack [27, 29].
@@
However, Designing a membership inference attack under partial feature knowledge introduces several challenges. We next introduce the challenges and our solutions one by one.
@@
Challenge 1: Incomplete inputs cannot be directly fed into the neural network, which makes it hard to get feedback from the model. The first step is therefore to determine how to fill in the missing features. Several studies have demonstrated that model inversion attacks can successfully recover input features[30, 31, 32, 33, 34, 35]. These attacks can be broadly categorized into two types. The first type aims to reconstruct representative features of an entire class, typically in scenarios where all samples from that class correspond to the same entity [30, 31, 32], and these attacks are generally conducted without prior knowledge of partial features. For example, in a facial recognition system, Fredrikson et al. [36] introduced the Face‑Rec attack, which reconstructs a human face using the target person’s label. The label, representing the target identity in the recognition system, serves as a guiding signal to steer the reconstruction toward the correct person. However, such attacks are not directly applicable to our setting, since in our case a class label does not uniquely correspond to a single entity. The second type of attack, also known as attribute inference attack, attempts to recover unknown private attributes based on the known features of a sample[34, 33, 35]. Yet, most existing works in this category assume that the attacker has prior knowledge of the possible values of the missing features, which is unrealistic in practice.
@@
Solution to Challenge 1: We adopt a reconstruction strategy similar to that of Fredrikson et al. [36], leveraging the model’s inherent memorization of its training data. Specifically, instead of using labels, we treat the known features as anchors and apply backpropagation to optimize the unknown features, aiming to minimize the sample’s loss value. Even when the known features originate from a training member, our goal is not to exactly reconstruct the original sample, but to recover what the model “remembers” , that is, the internal impression it has formed during training, typically corresponding to a point near a local minimum of the loss function. According to our observations, when the known features come from a member sample, the reconstruction tends to better align with the true data distribution, as the model has already formed internal representations based on the complete training inputs. In contrast, if the known feature combination is not a member, the optimization may converge to a random local minimum, resulting in a reconstruction that deviates more from the training distribution.
Challenge 2: It is difficult to distinguish members from non‑members. Because multiple samples may share the same known features, the reconstruction often becomes a blended representation influenced by several candidates, rather than a faithful replica of any individual one. As a result, the reconstructed sample differs from the original, making exact‑match membership inference attack no longer applicable. Moreover, without access to the full ground‑truth features, it is difficult to quantify the distance between the reconstructed and original samples, which also hinders the applicability of range‑match membership inference attacks.
@@
Solution to Challenge 2: We shift our focus from exact sample matching to assessing how well the reconstructed samples align with the underlying data distribution. Specifically, we employ anomaly detection techniques to identify reconstructed samples that deviate from the true data distribution, enabling us to distinguish between member and non-member feature combinations.
We provide a preliminary validation of the design intuition in Section V-B.

### v2

@@ Abstract
Machine learning models are vulnerable to membership inference attack, which can be used to determine whether a given sample appears in the training data. Most existing methods assume the attacker has full access to the features of the target sample. This assumption, however, does not hold in many real-world scenarios where only partial features are available, thereby limiting the applicability of these methods. In this work, we introduce Partial Feature Membership Inference (PFMI), a scenario where the adversary observes only partial features of each sample and aims to infer whether this observed subset was present in the training set. To address this problem, we propose MRAD (Memory-guided Reconstruction and Anomaly Detection), a two-stage attack framework that works in both white-box and black-box settings. In the first stage, MRAD leverages the latent memory of the target model to reconstruct the unknown features of the sample. We observe that when the known features are absent from the training set, the reconstructed sample deviates significantly from the true data distribution. Consequently, in the second stage, we use anomaly detection algorithms to measure the deviation between the reconstructed sample and the training data distribution, thereby determining whether the known features belong to a member of the training set. Empirical results demonstrate that MRAD is effective across various datasets, and maintains compatibility with off-the-shelf anomaly detection techniques. For example, on STL-10, our attack exceeds an AUC of around 0.75 even with 60% of the missing features.
@@ I Introduction
Machine learning has been widely applied across various domains, such as medical diagnosis[bilionis2025disparate], traffic analysis[hui2021trajnet], and autonomous driving[chen2024end]. However, current machine learning models remain vulnerable to privacy inference attacks. Adversaries can exploit interactions with the target model to infer privacy information from its training data, posing significant threats to user confidentiality and data privacy. A well-known example of such attacks is the Membership Inference Attack (MIA)[shokri2017membership, carlini2022membership], which can be used to determine whether certain samples were included in the training set of a target model. This type of attack highlights the risk of information leakage during inference, making the deployment of machine-learning services more challenging.
Membership inference attack has been shown to compromise the privacy of various machine learning models[fu2025unlocking, hu2025unveiling, galichin2025glira, chi2024shadow, wang2023link, zarifzadeh2024low]. As the majority of existing studies focus on classification models, our work follows the same line of research. Most membership inference attacks infer membership by computing attack signals, such as the sample loss, based on feedback from the target model together with prior knowledge of the model structure and training data[song2021systematic, ye2022enhanced, liu2023membership]. Under weaker attacker assumptions, some attacks are designed for settings with limited observability, where the attacker cannot obtain full model outputs[peng2024oslo, li2021membership] or lacks knowledge of the training data distribution and the target model details[liu2024gradient]. These settings further expand the practical scope of membership inference attacks. However, all of these methods share the same goal: determining whether a given sample exactly matches one of the training samples. Recently, Tao et al.[tao2025range] introduced range-based MIA as a new type of membership inference attack that breaks away from this exact-match paradigm. Instead of identifying whether a specific sample appears in the training set, their method aims to infer whether the training data contains any sample whose features fall within a specified range, highlighting that membership inference can extend beyond traditional exact-match assumptions.
@@
Inspired by Tao et al.[tao2025range], we pose a question: How much private information can membership inference attack extract when some features are missing? This setting aligns more closely with practical scenarios, as privacy constraints or instability in data sources may prevent an attacker from obtaining complete feature information for the target sample, leaving only partially observed tabular or image data available. To illustrate this, consider an attacker attempting to determine whether an individual, Alice, has a particular disease by querying a diagnostic system deployed in a hospital she frequently visits. If Alice is indeed diagnosed with the disease, it is likely that her data is included in the training set. However, due to the difficulty of acquiring complete information, the attacker may only know a subset of her attributes, such as age, gender, and race. Applying a standard membership inference attack in this setting would require exhaustively enumerating all possible combinations of the unknown features, which is computationally infeasible and inefficient.
@@
As Figure 1 shows, in this paper, we consider a setting where the attacker has access to only a subset of a sample’s features and aims to infer whether this partial feature combination was present in the training set of a target model. We refer to this type of attack as Partial Feature Membership Inference (PFMI). In this setting, the observed partial features is termed known features, while the remaining unobserved ones are referred to as unknown features. The property existence attack, proposed as a generalization of membership inference [chaudhari2023snap], shares certain similarities with PFMI in that both seek to infer the presence of specific feature combinations. However, in property existence attack, the adversary can directly access complete samples containing the target feature combination [chaudhari2023snap, wang2024property], whereas PFMI operates under stricter constraints by relying solely on partial observations. Although the attacker’s limited access to only partial feature information makes the inference task more challenging, this constraint also reflects a more realistic threat model. An adversary equipped with PFMI capabilities can still launch effective privacy inference attacks even with incomplete information. This significantly amplifies the risks faced by existing privacy protection mechanisms, which often assume that attackers have access to complete samples.
However, designing such attacks faces two key challenges:
  • Challenge 1: Traditional membership inference attacks rely on model feedback[carlini2022membership, zarifzadeh2024low], but incomplete samples cannot be directly input into the model to obtain meaningful responses.
  • Challenge 2: A new membership evaluation criterion is required to determine feature-level membership, as existing methods are no longer applicable in this setting.
@@
To tackle these problems, we propose MRAD—a two-stage attack framework that combines Memory-guided Reconstruction with Anomaly Detection. Specifically, to address the first challenge, the first stage uses the known features as anchors to recover the unknown features based on the representations learned by the model. Based on our observations, reconstructed samples with non-member features often deviate significantly from the original distribution, resulting in anomalous data. We leverage anomaly detection techniques to identify such deviations, thus determining whether the observed partial feature combination exists in the training set, thereby addressing the second challenge. Further details will be provided in Section IV-A. Our main contributions can be summarized as follows.
@@
  • We investigate a more realistic setting for membership inference attack, in which the adversary has access only to partial features of the target sample. We refer to this scenario as PFMI, and we provide a formal definition of the attack framework under this constraint.
@@
  • We propose a two-stage attack framework designed for this new inference scenario, which is applicable to both white-box and black-box settings. The framework is modular and can be integrated with a variety of off-the-shelf anomaly detection techniques.
@@
  • We extensively evaluate our attack framework on both image and tabular datasets, demonstrating its compatibility with various anomaly detection methods. For the STL-10 dataset, our method still exceeds an AUC of 0.75 even when 60% of the features are missing.
@@
  • We show how hyperparameter settings and the importance of observed features affect attack performance by experiments. We also present a case study to demonstrate how our attack can lead to privacy leakage in a real world scenario.
@@
The remainder of this paper is organized as follows. Section II reviews related work on membership inference and anomaly detection. Section III introduces the PFMI setting, including the attack objective, adversary capability and attack definition. Section IV presents the technical details of the proposed two-stage attack framework. Section V presents experimental results, including studies that validate design intuition, followed by a comprehensive evaluation of the attack performance in various settings. Section VI analyzes the relationship between PFMI and other inference attacks, and outlines directions for future work. Section VII concludes the paper.
@@ II-A Membership Inference Attack
Membership inference attack against machine learning models was first introduced by Shokri et al.[shokri2017membership]. A binary classifier was trained as the attack model using prediction vectors produced by shadow models. To launch the attack, the target model’s prediction on a given input is provided to the attack model, which decides whether the input was included in the training data. Most follow-up studies adopt a similar setting, where the adversary has access to rich information, such as the full model outputs and the training data distribution. These methods typically rely on the observation that member samples tend to yield lower loss values than non-members [song2021systematic, sablayrolles2019white, carlini2022membership]. Song et al.[song2021systematic] predict membership with a modified loss function. Ye et al.[ye2022enhanced] proposed a membership inference attack that achieves an arbitrary false positive rate by using a non-member set to calibrate the decision threshold. Carlini et al. [carlini2022membership] trained multiple shadow models to examine the loss distributions of samples when they are in and out of the training set, further improving the attack performance under low false positive rates. Instead of using loss values for membership inference, Zarifzadeh et al.[zarifzadeh2024low] computed a likelihood ratio based on the output confidences and used it as the inference metric.
@@
Another line of work considers membership inference under restricted attacker capabilities. In the label-only setting, the adversary can only observe the predicted label rather than the full confidence vector [li2021membership, choquette2021label, peng2024oslo]. Li et al. [li2021membership] leveraged the fact that member samples are often farther from the decision boundary, while Peng et al. [peng2024oslo] crafted minimally perturbed adversarial examples using a shadow model and inferred membership based on whether the perturbed samples were misclassified by the target model. As the method proposed by Peng et al.[peng2024oslo] relies solely on the final prediction of the target model, it requires only a single query. Another setting is blind membership inference, in which the adversary has access to the full prediction vector of the target sample but lacks knowledge of the true label. Hui et al. [hui2021practical] proposed a blind black-box attack that performs differential comparisons between samples to infer membership without ground-truth guidance. In addition, Liu et al.[liu2024gradient] demonstrate that membership inference remains feasible even when the adversary lacks prior knowledge of both the target model and the training data.
@@
Recently, Tao et al.[tao2025range] proposed a novel form of membership inference attack that shifts the focus from individual samples to feature regions. Rather than determining whether a specific sample was included in the training set, their method aims to infer whether any data point within a given range appears in the training data. To perform the attack, they first sample instances from the specified range, apply standard membership inference techniques to assign scores to these samples, and then aggregate the scores to reach a final inference decision.
@@ II-B Anomaly Detection
Anomalous data refers to samples that deviate from the expected or learned data distribution. Two common categories of such anomalies are concept drift and out-of-distribution (OOD) samples. Concept drift occurs when the test-time data distribution shifts from the training distribution, leading to degraded model performance [wares2019data]. Yang et al. [yang2021cade] detect such drift by modeling input data with autoencoders and computing Median Absolute Deviation(MAD) values between the current and original distributions. Wan et al. [wan2024online] proposed MCD-DD, which uses a dynamically updated encoder and Mahalanobis distance-based criteria to detect distributional changes in time series. OOD detection has also received significant attention, as the presence of OOD samples can bias model behavior. Park et al. [park2023nearest] improved OOD detection by reducing overconfidence in OOD regions. Reiss et al. [reiss2023mean] designed a novel contrastive loss function that enhances contrastive learning for OOD detection. Liu et al. [liu2025detecting] exploited the clustering effect in the penultimate layer of neural networks to develop a post-hoc OOD detection method.
@@ III-A Attack Model
Attack Objective: In this work, the goal of the attacker is similar to that of traditional MIA, but with a key constraint: the attacker only has access to partial features of the target sample. In this paper, for the sake of generality, we define partial features as a subset of the full feature vector, where the known features can appear at arbitrary positions. That is, the known and unkonwn features are not required to form a contiguous block or follow any fixed pattern. This flexible definition better captures real-world scenarios where feature availability is influenced by privacy constraints or unstable data sources. The objective is to determine whether this particular feature combination exists in the training set of the target model. We refer to such feature combinations as member features if they appear in the training set, and non-member features otherwise. Apparently, any combination drawn from the training set is considered as member features.
Adversary Capability: This work considers both white-box and black-box attack settings. In the white-box setting, the attacker has full access to the target model, including its outputs, training algorithm, internal parameters, and gradients. In contrast, under the black-box setting, the attacker can only obtain the full prediction vector output by the target model. In addition, although the specific target sample may have missing features due to privacy concerns or incomplete collection, we assume that the attacker has access to fully observed auxiliary data drawn from the same underlying distribution as the target model training data. These auxiliary samples are used to support the attack, for example by training proxy models or estimating relevant statistics, which is a common setting in prior membership inference attack settings[shokri2017membership, song2021systematic, ye2022enhanced, zarifzadeh2024low]. This is feasible in practice, as attackers may obtain partial data through social engineering or cyberattacks. However, the attacker is unable to obtain any complete sample that contains the exact feature combination to be inferred.
@@ III-B Attack Definition
Before introducing our proposed attack definition, we first revisit the standard definition of membership inference attack, as formalized in prior work[yeom2018privacy, zarifzadeh2024low] using an indistinguishability game.
@@ IV-A Design Intuition
In designing our attack, we build on a well‑established observation: overfitted models memorize their training data, causing training members to occupy regions of the input space associated with low loss values[shokri2017membership, salem2023sok, hu2022membership]. Specifically, for a supervised machine learning algorithm $\mathcal{T}$, the training process aims to learn a model that can effectively distinguish between samples from different classes. Since it is impractical to access the full data distribution, a finite subset of labeled examples $\mathcal{D}_{train}$, the training set, is typically sampled to represent it. A function $f_{\theta}$, parameterized by $\theta$, is then learned to map each input $x$ from $\mathcal{D}_{train}$ to its corresponding label $k$ by minimizing a loss function $\mathcal{L}_{\theta}(x,k)$. Because the optimization only considers the training set, the model tends to assign lower loss values to training samples compared to unseen ones. This results in loss valleys around training points in the input space, a phenomenon that forms the basis of many membership inference attack [yeom2018privacy, hu2022membership].
@@
However, as discussed in Section I, Designing a membership inference attack under partial feature knowledge introduces several challenges. We next provide a detailed elaboration of these challenges and their corresponding solutions.
@@
Challenge 1: Incomplete inputs cannot be fed into the neural network directly, which makes it hard to get feedback from the model. The first step is therefore to determine how to fill in the missing features. Several studies have demonstrated that model inversion attacks can successfully recover input features[dibbo2023model, kahla2022label, an2022mirror, annamalai2024linear, mehnaz2022your, kabir2025disparate]. These attacks can be broadly categorized into two types. The first type aims to reconstruct representative features of an entire class, typically in scenarios where all samples from that class correspond to the same entity [dibbo2023model, kahla2022label, an2022mirror], and these attacks are generally conducted without prior knowledge of partial features. For example, in a facial recognition system, Fredrikson et al. [fredrikson2015model] introduced the Face‑Rec attack, which reconstructs a human face using the target person’s label. The label, representing the target identity in the recognition system, serves as a guiding signal to steer the reconstruction toward the correct person. However, such attacks are not directly applicable to our setting, since in our case a class label does not uniquely correspond to a single entity. The second type of attack, also known as attribute inference attack, attempts to recover unknown private attributes based on the known features of a sample[mehnaz2022your, annamalai2024linear, kabir2025disparate]. Yet, most existing works in this category assume that the attacker has prior knowledge of the possible values of the missing features, which is unrealistic in practice.
@@
Solution to Challenge 1: We adopt a reconstruction strategy similar to that of Fredrikson et al. [fredrikson2015model], leveraging the model’s inherent memorization of its training data. Specifically, instead of using labels, we treat the known features as anchors and apply backpropagation to optimize the unknown features, aiming to minimize the sample’s loss value. Our goal is to recover what the model “remembers” , that is, the internal impression it has formed during training, typically corresponding to a point near a local minimum of the loss function. Due to model memorization, the differences present in reconstructed samples serve as crucial evidence for inferring membership.
Challenge 2: Membership inference becomes more challenging when feature representations are ambiguous. Multiple samples may share the same known features, instead of reproducing the true underlying sample, the reconstruction often collapses into a blended or averaged representation shaped by several plausible candidates. This prevents the model from faithfully capturing the characteristics of any single instance, causing the reconstructed result to deviate noticeably from the original. As a consequence, membership inference methods that rely on detecting exact matches become fundamentally ineffective.
Moreover, since the attacker never observes a fully specified instance with the target feature combination, there is no principled way to characterize the feasible space of the unknown attributes. Without a well-defined feature range, it is not possible to establish meaningful boundaries for sampling, effectively ruling out the use of range-based attack proposed by Tao et al.[tao2025range] in this setting.
@@
Solution to Challenge 2: We shift our focus from exact sample matching to assessing how well the reconstructed samples align with the underlying data distribution. According to our observations, when the known features come from a member sample, the reconstruction tends to better align with the true data distribution, as the model has already formed internal representations based on the complete training inputs. In contrast, if the known feature combination is not a member, the optimization may converge to a random local minimum, resulting in a reconstruction that deviates more from the training distribution. Based on this observation, we employ anomaly detection techniques to identify reconstructed samples that deviate from the true data distribution, enabling us to distinguish between member and non-member feature combinations. The experiment results in Section V-B confirm the validity of our observations.